# AcousTools: A ‘Full-Stack’, Python-Based, Acoustic Holography Library

### v6

@@ 4.2 Propagators
Because these propagators share the same structure - both are matrices where the complex pressure at a point is obtained by the matrix-vector product of the propagator and the transducer activations - these can be swapped without requiring additional changes to the rest of the code. This allows a user to compare the two contexts with minimal changes to their code. For some solver (e.g., IB[IB]), the propagator is the only component that needs to be interchanged between the piston model and BEM contexts.
@@ 4.3 Solvers
AcousTools provides a number of solvers, giving further flexibility depending on the context of an application. High-speed solvers, such as multipoint naive and GS-PAT, offer quick computation[GSPAT], while slower solvers, such as WGS, can generate focal points with small variance of pressure [TGS]. As discussed, the choice of which solver to use is a critical stage in the development process, and a wide variety that can be easily compared enables a developer to make an informed choice.
@@
AcousTools can leverage implicit-gradient descent algorithms to optimise more complex objectives [Diff-PAT, Diff-PAT-Expanded], which can go beyond what projective algorithms can do. Implicit gradient-descent methods can minimise Gor’kov potential directly, produce fields with specific metrics at specific points, and even combine multiple objectives simultaneously. Although these methods are often slower than others [Diff-PAT] and may not be suitable for all cases, their ability to use custom objectives provides greater flexibility for the developer.
@@
This results in the sound field shown in Fig. 4, using the Adam optimiser[Adam], a learning rate of $0.01$, 10001000 iterations and $\lambda=1.32\times 10^{10}$ which ensures each component has a similar magnitude. One point has a minima in the Gor’kov potential corresponding to a low pressure region sandwiched between two high pressure regions and the other point has a maxima of Gor’kov potential corresponding to a region of high pressure, as defined in the objective. This could be useful for multi-modal applications where a focal point is used for haptic applications while the minima of Gor’kov potential corresponds to a levitation trap that can hold a small bead [AcousticElements].
Projective solvers, such as IBP and GS-PAT, could not fully create the field in this situation, as they produce high pressure regions at their targets. These can be converted to traps through levitation signatures [AcousticElements], but such signatures convert all focal points instead of only one. As seen in Fig. 4, a focal point has a local minima just below the high pressure regions, and so a similar result to Fig. 4 could be produced by focusing slightly above the target trap position. However, using this method may lead to both sub-optimal solutions with smaller forces and is harder to customise the relative strengths of the trap and focus. Therefore, the implicit-gradient solver is an additional tool for a developer, and combining this with the other methods provides a full toolbox to use.
@@ 4.4 Analysis
Computing Gor’kov potential here uses analytical gradients of the piston model instead of using finite differences and makes it significantly faster compared to the standard method using finite differences [AcousticElements, HighSpeedBEM]. The speed-up between the two methods is significant, the analytical method is able to compute 4232 solutions per second compared to 1922 when using finite differences; these values were computed over 3000 random independent points, computed sequentially. A similar analytical solution is provided using the analytical gradient of BEM. For both propagators, an analytical method for computing the gradient of Gor’kov potential is also provided for computing force. A summary of these metrics can be seen in Fig. 5, showing that all these metrics from pressure to stiffness. In this case, the propagator used is the piston model and metrics use the analytical gradients where possible, similar visulisations could be created using BEM by simply swapping the propagators as discussed above.
For a developer, the ability to select the appropriate metric for a given situation ensures that the acoustic behaviour can either be quantified after the fact to verify correct performance [Ultraino] or included in the solver to allow explicit optimisation of these quantities [VolDisplaySingleParticle, Diff-PAT, AcousticElements]. Once again, a versatile selection leads to more versatile applications for developers.
@@ 4.5 Hardware
The final stage of the full-stack is Hardware, sending the transducer activations to a real world device. As discussed, this is difficult for a software package to be fully general purpose as every device may have a different communication protocol with little standardisation. Therefore, AcousTools uses OpenMPD’s board drivers [OpenMPD] to drive the 16x16 transducer arrays used in that paper. This is because these arrays are common across a number of applications, from levitation[VolDisplaySingleParticle, GSPAT], haptics [HapticsMPSTM, ToBEMorNot] to fabrication[AcoustoFab], have been the target of many phase retrieval methods [Diff-PAT, GSPAT, VolDisplaySingleParticle] and propagation methods [AcousticElements, HighSpeedBEM], and are based on an open source platform[Ultraino, OpenMPD]. Sending the activations is simple with AcousTools. All the developer must do is connect to the device and send the computed activations:
@@
Figure 6: Examples of real-world levitation prodiced with AcousTools using two opposing transducer arrays. a) 8 points forming a static cube. b) A circle rendered using one particle moving at high speed, using persistence of vision effects to produce the illusion of a single shape. Note in the high speed case an acoustically-transparent mesh covers the bottom array to ensure particles are not dropped into the transducers but this does not affect acoustic behaviour.
@@ 4.6 Application: Droplet-Solid Levitation
The acoustic levitation of droplets is desirable from applications from fabrication[Levitate, AcoustoFab] to mid-air chemistry[AcousticChemistry] but is more challenging than solid particle levitation because too strong of a force applied to the droplet may cause it to buckle and burst [Droplets, HighSpeedBEM]. This can be mitigated by optimising the trap to have a reduced Gor’kov potential applying a controlled force to the droplet. On the other hand, the solid particle is best trapped by minimising Gor’kov potential and therefore applying high, converging forces. If traps of the same strength are used for both traps, the droplet would be at risk of bursting if both are strong or the solid bead would have sub-optimal forces if the traps are weaker, meaning if the particle is moved then it will be at risk of being dropped[VolDisplaySingleParticle]. Careful control of the forces gives a best-of-both-worlds situation, with the forces being as strong as possible for both traps without harming reliability [Droplets].
@@
Here, $\boldsymbol{E}_{1}$ and $\boldsymbol{E}_{2}$ are the BEM propagators that model the scattering from a flat reflector to each of the target points. The resulting levitation as well as the simulated pressure and Gor’kov potential are shown in Fig. 7. This was conducted using a $U_{target}=-1\times 10^{-7}J$ for the droplet trap and $\lambda=7\times 10^{8}$. A learning rate of $1$ was used and the optimisation was performed over $1000$ steps with the Adam optimiser[Adam]. The droplet trap has a resulting Gor’kov potential, $U=-9.98\times 10^{-8}J$, $99.8\%$ of the desired target.
@@ 5 Discussion
Developers from a wide variety of fields can take AcousTools and deploy it in their contexts. Python as a language is among the easiest languages to learn [pythonEase] and is one of the most used languages [PythonUse] meaning there is a very large number of people who would be able to take AcousTools and integrate acoustic holography into their own domain. A haptician could compare the effects of different transducer board or different propagators on immersion, someone interested in levitation could optimise the shapes that they render to give higher visual quality or a chemist could control small amounts of liquid drops for non-contact mid-air reactions while maintaining specific forces. AcousTools’ use of PyTorch not only provides the ability to integrate implicit-gradient solvers as discussed above but also allows users to deploy a machine learning solutions to their work. Novel solutions that have already been shown to be helpful for various acoustic domains [AcousNet, StableLev]. Thanks to AcousTools, possible applications in acoustic holography are not limited by existing software and now depend on developers alone.

### v7

@@ 4.2 Propagators
Because these propagators share the same structure - both are matrices where the complex pressure at a point is obtained by the matrix-vector product of the propagator and the transducer activations - these can be swapped without requiring additional changes to the rest of the code. This allows a user to compare the two contexts with minimal changes to their code. For some solver (e.g., IB [31]), the propagator is the only component that needs to be interchanged between the piston model and BEM contexts.
@@
These propagators all assume that the medium is homogeneous - for other applications that this assumption is not true, such as in-vivo applications, other applications may need more complex propagators. Critically though, if they take the same form as the existing models, they can then be used with the rest of AcousTools as with any other propagation matrix.
@@ 4.3 Solvers
AcousTools provides a number of solvers, giving further flexibility depending on the context of an application. High-speed solvers, such as multipoint naive and GS-PAT, offer quick computation [41], while slower solvers, such as WGS, can generate focal points with small variance of pressure [6]. As discussed, the choice of which solver to use is a critical stage in the development process, and a wide variety that can be easily compared enables a developer to make an informed choice.
@@
AcousTools can leverage implicit-gradient descent algorithms to optimise more complex objectives [11, 12], which can go beyond what projective algorithms can do. Implicit gradient-descent methods can minimise Gor’kov potential directly, produce fields with specific metrics at specific points, and even combine multiple objectives simultaneously. Although these methods are often slower than others [11] and may not be suitable for all cases, their ability to use custom objectives provides greater flexibility for the developer.
@@
This results in the sound field shown in Fig. 4, using the Adam optimiser [24], a learning rate of $0.01$, 10001000 iterations and $\lambda=1.32\times 10^{10}$ which ensures each component has a similar magnitude. One point has a minima in the Gor’kov potential corresponding to a low pressure region sandwiched between two high pressure regions and the other point has a maxima of Gor’kov potential corresponding to a region of high pressure, as defined in the objective. This could be useful for multi-modal applications where a focal point is used for haptic applications while the minima of Gor’kov potential corresponds to a levitation trap that can hold a small bead [32].
Projective solvers, such as IBP and GS-PAT, could not fully create the field in this situation, as they produce high pressure regions at their targets. These can be converted to traps through levitation signatures [32], but such signatures convert all focal points instead of only one. As seen in Fig. 4, a focal point has a local minima just below the high pressure regions, and so a similar result to Fig. 4 could be produced by focusing slightly above the target trap position. However, using this method may lead to both sub-optimal solutions with smaller forces and is harder to customise the relative strengths of the trap and focus. Therefore, the implicit-gradient solver is an additional tool for a developer, and combining this with the other methods provides a full toolbox to use.
@@
The main other category of phase-retrieval algorithms is ML based methods. As discussed, these use large networks to predict the phases that create some target image in the focal plane. These are often specialist networks and are not directly implemented in AcousTools but because it is built upon PyTorch, every element of AcousTools is able to be used with the automatic differentiation engine, meaning ML components can be designed, trained and deployed using it. Because of this, AcousTools can facilitate the growth of this corner of the acoustics world - providing a combined and unified platform for research.
@@ 4.4 Analysis
Computing Gor’kov potential here uses analytical gradients of the piston model instead of using finite differences and makes it significantly faster compared to the standard method using finite differences [32, 18]. The speed-up between the two methods is significant, the analytical method is able to compute 4232 solutions per second compared to 1922 when using finite differences; these values were computed over 3000 random independent points, computed sequentially. A similar analytical solution is provided using the analytical gradient of BEM. For both propagators, an analytical method for computing the gradient of Gor’kov potential is also provided for computing force. A summary of these metrics can be seen in Fig. 5, showing that all these metrics from pressure to stiffness. In this case, the propagator used is the piston model and metrics use the analytical gradients where possible, similar visulisations could be created using BEM by simply swapping the propagators as discussed above.
For a developer, the ability to select the appropriate metric for a given situation ensures that the acoustic behaviour can either be quantified after the fact to verify correct performance [30] or included in the solver to allow explicit optimisation of these quantities [19, 11, 32]. Once again, a versatile selection leads to more versatile applications for developers.
@@ 4.5 Hardware
The final stage of the full-stack is Hardware, sending the transducer activations to a real world device. As discussed, this is difficult for a software package to be fully general purpose as every device may have a different communication protocol with little standardisation. Therefore, AcousTools uses OpenMPD’s board drivers [36] to drive the 16x16 transducer arrays used in that paper. This is because these arrays are common across a number of applications, from levitation [19, 41], haptics [45, 37] to fabrication [5], have been the target of many phase retrieval methods [11, 41, 19] and propagation methods [32, 18], and are based on an open source platform [30, 36]. Sending the activations is simple with AcousTools. All the developer must do is connect to the device and send the computed activations:
@@
Figure 6: Examples of real-world levitation prodiced with AcousTools using two opposing transducer arrays. a) 8 points forming a static cube. b) A circle rendered using one particle moving at high speed, using persistence of vision effects to produce the illusion of a single shape.
@@ 4.6 Application: Droplet-Solid Levitation
The acoustic levitation of droplets is desirable from applications from fabrication [1, 5] to mid-air chemistry [42] but is more challenging than solid particle levitation because too strong of a force applied to the droplet may cause it to buckle and burst [26, 18]. This can be mitigated by optimising the trap to have a reduced Gor’kov potential applying a controlled force to the droplet. On the other hand, the solid particle is best trapped by minimising Gor’kov potential and therefore applying high, converging forces. If traps of the same strength are used for both traps, the droplet would be at risk of bursting if both are strong or the solid bead would have sub-optimal forces if the traps are weaker, meaning if the particle is moved then it will be at risk of being dropped [19]. Careful control of the forces gives a best-of-both-worlds situation, with the forces being as strong as possible for both traps without harming reliability [26].
@@
Here, $\boldsymbol{E}_{1}$ and $\boldsymbol{E}_{2}$ are the BEM propagators that model the scattering from a flat reflector to each of the target points. The resulting levitation as well as the simulated pressure and Gor’kov potential are shown in Fig. 7. This was conducted using a $U_{target}=-1\times 10^{-7}J$ for the droplet trap and $\lambda=7\times 10^{8}$. A learning rate of $1$ was used and the optimisation was performed over $1000$ steps with the Adam optimiser [24]. The droplet trap has a resulting Gor’kov potential, $U=-9.98\times 10^{-8}J$, $99.8\%$ of the desired target.
@@ 5 Discussion
Developers from a wide variety of fields can take AcousTools and deploy it in their contexts. Python as a language is among the easiest languages to learn [47] and is one of the most used languages [25] meaning there is a very large number of people who would be able to take AcousTools and integrate acoustic holography into their own domain. A haptician could compare the effects of different transducer board or different propagators on immersion, someone interested in levitation could optimise the shapes that they render to give higher visual quality or a chemist could control small amounts of liquid drops for non-contact mid-air reactions while maintaining specific forces. AcousTools’ use of PyTorch not only provides the ability to integrate implicit-gradient solvers as discussed above but also allows users to deploy a machine learning solutions to their work. Novel solutions that have already been shown to be helpful for various acoustic domains [51, 13]. Thanks to AcousTools, possible applications in acoustic holography are not limited by existing software and now depend on developers alone.